Telescope Bootstrap Startup: 11 September 2012 UT

Observer: JHill (LBTO)
Telescope Operator: DHuerta (LBTO)
Software Support: CBiddick (Tucson)
AO Support: DMiller (Tucson)
Telescope Support: KNewton and JLittle(LBTO)
Instrument Support: JMorris (LBTO)

Summary

This night was lost to weather. There were heavy thunderstorms in the afternoon, and then light rain starting around 4UT. The sky improved to partly cloudy after midnight, but the dome never dried off.

Preparations

M3 selector

Torsten gave us a patch to OSS (v 2.25.2) with mutexes that fixed the M3 selector problem - Issue 4194. This was tested here. The same patch has configureable timeouts to solve Issue 4119 which we tested below.

00:35 stopped and restarted OSS to get the patched version

00:36 Authorize LBTI on both sides (M3s starting at LFBG/RFBG) - success, both selectors move correctly in parallel.

00:43 Authorize LUCI on both sides (M3s starting at LBTI) - success, both selectors move correctly in parallel.

00:46 Authorize LBTI on both sides (M3s starting at LFBG/RFBG) - success, both selectors move correctly in parallel. It is interesting to note that the Authorize command returns success a few seconds before the M3 tip-tilt finishes moving. (I believe that is a feature rather than a bug.)

Other

00:40 Daytime AO testing concludes.

01:20 Some water dripped on the DX primary during the afternoon storm. We were able to blow it off with dry compressed air. This was a different leak in a different place than the one that made the rusty water marks in August.

We inspected the scratches on the SX primary and found them to be fairly small. See Issue 4191 for details.

Details

03:40 We were briefly encouraged that the humidity was only 70% outside. Then we noticed that it was raining.

LBTI AO tests

04:00 Vidhya is testing the LBTI AO software on the left side.

05:xx David reports some SERDES errors. That turned out to be because the treehouse cooling was still off.

06:xx We had two instances where we had to reinitialize the SX M2 hexapod. We don't know why.

07:15 End of Vidhya's testing for tonight. She still has a couple commands that don't work through the ICE interface.

07:20 The rain has stopped, and the sky is clearing off. Humidity is fluctuating from 70 to 83%. More clouds can be seen upstream to the south in the satellite image.

Additional OSS software tests

added these to oss.conf so we can test Torsten's new configurable timeouts
oss.terc.selector.timeout int 300                 # timeout for M3 selector motions (sec) (300) implemented 20120911
oss.terc.tiptilt.timeout int 60                   # timeout for M3 tip tilt piston motions (sec) (60) implemented 20120911

07:26 stop and restart OSS (this is not needed for this test, it reads the file for every move)

07:28 Authorize LUCI on both sides - M3 moved, but failed because SX shell not set.

07:30 set SX shell, collimate SX M3

07:32 authorize LBTI on both sides - M3 moved,

07:37 authorize IRTC/LUCI

Tue Sep 11 07:38:46.619 2012 TCS authorize alert for left IRTC at bentGregorianFront focus and right LUCIFER at bentGregorianFront focus: adsecarb:Illegal command for state Ready (Load shape)

07:39 send $$$ on SX to reboot M3 UMAC (I need something to make a command fail if I want to see the timeout.)

07:40 Press "Collimate" on PSF SX M3 GUI - and the mirror moves to the correct position even though OSS M3 GUI is saying NOT Homed.

07:41 De-init SX M3

07:42 Press "Collimate" on PSF SX M3 GUI - it still works

07:43 Press "Init" on both M3s within 1 sec of each other (SX first, then DX). DX succeed in homing the selector, but SX looks like it failed to move the selector and then got a false positive limit indication; alternately the positive limit was correct and the position feedback was wrong. I'm not sure which.

07:48 send $$$ on SX to reboot M3 UMAC

07:49 Init SX M3 alone

07:52 Init DX M3 alone

07:57 Set tiptilt timeout to an unreasonably short 5 sec.

We don't get timeouts on tip-tilt moves sent from OSS GUI or from PSF. The tcslog says:
Sep 11 07:58:26 tcs1 LBT_OSS: [D] waiting for the tertiary to start. Timeout for this move is 300 seconds.
OSS seems to know that I changed the file, but it doesn't seem to have digested my timeout in the file.

Aha . This is because Torsten previously put the timeouts higher up in the file where I didn't see them.

08:16:39 Now my very short 5 sec timeout works correctly.

08:18 Set tiptilt timeout to 60 sec

08:19 Authorize IRTC and LUCIFER (M3s were at LBTI) - OK

08:21 Authorize LBTI on both sides - OK

-- JohnHill - 11 Sep 2012
Topic revision: r2 - 11 Sep 2012, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback